OpenStudy (anonymous):

34 ÷ (14 − 5) × 2. Its 3 to the power of 4 not 34.

OpenStudy (decentnabeel):

@KAKES1967 you need to just answer

OpenStudy (anonymous):

huh?

OpenStudy (decentnabeel):

answer is 18

OpenStudy (decentnabeel):

fist solve the parenthesis

OpenStudy (decentnabeel):

then apply DMAS rule

OpenStudy (mertsj):

\[3^4\div(14-5)\times 2\] \[81\div 9 \times 2\] \[9 \times 2 = 18\]

OpenStudy (anonymous):

Thank you both c:

OpenStudy (decentnabeel):

No problem :) @KAKES1967